# From Pixels to Titles: Video Game Identification by Screenshots using Convolutional Neural Networks

### v2

@@ Abstract
This paper investigates video game identification through single screenshots, utilizing five convolutional neural network (CNN) architectures (MobileNet, DenseNet, EfficientNetB0, EfficientNetB2, and EfficientNetB3) across 22 home console systems, spanning from Atari 2600 to PlayStation 5, totalling 8,796 games and 170,881 screenshots. Confirming the hypothesis, CNNs autonomously extract image features, enabling the identification of game titles from screenshots without additional features. Using ImageNet pre-trained weights as initial weights, EfficientNetB3 achieves the highest average accuracy (74.51%), while DenseNet169 excels in 14 of the 22 systems. Employing alternative initial weights trained in an arcade screenshots dataset boosts accuracy for EfficientNetB2 and EfficientNetB3, with the latter reaching a peak accuracy of 76.36% and demonstrating reduced convergence epochs from 23.7 to 20.5 on average. Overall, the combination of optimal architecture and weights attains 77.67% accuracy, primarily led by EfficientNetB3 in 19 systems. These findings underscore the efficacy of CNNs in video game identification through screenshots.
keywords:
@@
video game identification, convolutional neural networks, transfer learning, single screenshot analysis, automated game recognition
1 Introduction
@@
Humans possess the remarkable ability to easily recognize their favorite video games or titles they have played frequently from a single screenshot. This proficiency is rooted in the presence of consistent visual elements, including sprites, heads-up displays (HUDs), and distinctive game scenarios. However, extending this capability to the automated identification of random video games from an extensive console library presents a formidable challenge, even for the most dedicated gamers. Therefore, the concept of automatically identifying video games from single screenshots holds immense appeal, not only for its technical complexity but also for its vast practical applications.
@@
Automated game identification could offer substantial benefits to various sectors within the gaming industry. Video game databases, search engines, and online platforms stand to gain significantly from this technology. By analyzing user-uploaded screenshots, these platforms can automatically generate metadata, including game titles, release dates, and developer information. Such automation would not only improve the accuracy of their game libraries but also enhance cataloging efficiency. Moreover, online streaming platforms could harness screenshot recognition to provide real-time information to viewers about the games being played during live streams, enhancing the overall viewer experience. This technology opens doors to further innovation within the gaming ecosystem, potentially influencing game recommendation systems and aiding game-related research.
@@
Most of the video games classification attempts so far aimed at genre classification. Augusto De Souza et al. [1] pioneering work classified game genre of gameplay videos. Their dataset comprises 700 gameplay videos spanning seven distinct game genres. In their research, they introduced novel descriptors known as Bossa Nova and BinBoost. The experimental outcomes demonstrated the effectiveness of their proposed approach, achieving an accuracy rate of 89.84%.
@@
While the majority of classification endeavors have typically focused on broader categories, there was a unique attempt to classify video games by their titles more than a decade ago. In this pioneering effort, Madani et al. [4] explored several fusion methods using a dataset containing 120,000 gameplay videos, with the objective of identifying game titles. Their approach integrated both audio and visual features to accurately pinpoint these specific game titles, ultimately achieving a F1-score of 0.82. Although their dataset is considerably large, they explored only a small number of games, identifying 30 distinct game titles. To the best of my knowledge, there has not been any other attempt to identify video games by their titles, especially in larger datasets with hundreds or thousands of titles, neither using video nor screenshots.
In the past decade, there has been a notable surge in the adoption of deep learning methods, as exemplified by the works of LeCun et al. [5], Goodfellow et al. [6], and Schmidhuber [7]. Among these methods, convolutional neural networks (CNNs) have played a pivotal role in driving advancements in automatic image classification, as initially demonstrated by Krizhevsky et al. [8]. CNNs represent a specific category of deep neural networks that find widespread use in the field of visual image analysis.
However, one of the inherent challenges associated with CNNs lies in their substantial appetite for annotated image samples, which are essential for estimating the millions of parameters required for network training. The process of annotating images can be both expensive and time-consuming, often presenting a significant bottleneck, particularly when dealing with problems characterized by limited available training data [9]. This limitation has, at times, hindered the widespread application of CNNs in scenarios where access to abundant training data is restricted.
Remarkably, Oquab et al. [9] addressed this predicament by unveiling a breakthrough approach. Their research demonstrated that image representations acquired through CNNs trained on extensive, annotated datasets could be judiciously leveraged for other visual recognition tasks, even in cases where only a limited amount of training data is available. This innovative method involved repurposing layers from a CNN model previously trained on a large dataset to compute mid-level image representations for a different dataset, yielding remarkable improvements in classification performance. This powerful technique, commonly referred to as transfer learning, has found successful application across various domains and scenarios, as exemplified by the works of Shin et al. [10] Huynh et al. [11] and Gopalakrishnan et al. [12].
@@
Recently, Jiang and Zheng [15] devised deep neural networks for the purpose of classifying game genres using either cover images or description text. Their dataset encompassed cover images and description texts sourced from a pool of 50,000 games, which they categorized into 15 distinct genres. In their approach, several pre-trained CNNs were fine-tuned for the cover image classification task. For the classification of description text, they employed Long Short-Term Memory (LSTM) networks and the Universal Sentence Encoder (USE). The image-based model yielded a highest accuracy rate of 31.4% when utilizing ResNet-50. They also achieved significant improvement in accuracy, up to 49.9%, by combining image and text features within a multi-modal model.
Both of these previous studies were limited to identifying game genres, and their accuracy was only moderate (below 50%). Until now, there has been no attempt to identify game titles using CNNs, which is a more challenging task than identifying only the game genres. In this paper, the primary focus revolves around the task of video game title classification based on single screenshots, utilizing CNN models. The hypothesis is that the inherent CNN capacity of automatically extracting relevant features from images is sufficient to identify video game titles from single screenshots in most scenarios, without relying on other features. To embark on this journey, a dataset encompassing 170,881 screenshots from 8,796 games of 22 popular home console systems was curated. The screenshots were sourced from the reputable Moby Games Database [16]. The proposed dataset spans a wide spectrum of gaming history, ranging from iconic consoles like the ‘Atari 2600’ of the second generation to the cutting-edge ‘PlayStation 5’ and ‘Xbox Series’ of the current generation, carefully selecting the most sold consoles from each generation between them. Its richness and diversity make it an ideal playground for our research.
To tackle this ambitious task, well-established CNN architectures were selected: MobileNet [17], DenseNet [18], and EfficientNet [19]. These architectures have consistently demonstrated outstanding performance in previous works with different kinds of images [20, 21], making them prime candidates for this game title classification task. The initial weights of these CNNs were first initialized with pre-trained weights from the ImageNet dataset [22], a widely adopted approach in transfer learning [9]. Subsequently, the pre-trained weights from another dataset of screenshots were employed to enhance both classification accuracy and reduce training times.
What sets this research apart is its pioneering spirit. To the best of my knowledge, this marks the first attempt to tackle the intriguing challenge of game title classification using CNNs. By pushing the boundaries of automated video game identification, the aim is to contribute valuable insights to game-related research and practical applications to the ever-evolving gaming industry.
The remainder of this paper is organized as follows. Section 2 presents the Moby Games Database and how the dataset was sourced from it. Section 3 shows the CNN architectures employed in this paper. Section 4 displays the computer simulations comparing the CNN architectures in the task of identifying the games from their screenshots, initialized with pre-trained weights from the ImageNet dataset. Section 5 demonstrates computer simulations using weights pre-trained in another screenshots dataset, comparing the accuracy and training epochs with those obtained with the ImageNet weights. Finally, the conclusions are drawn in Section 6.
2 The Dataset
@@
The Moby Games Database [16], as stated on their website, is an ambitious project with the primary goal of meticulously cataloging comprehensive information about electronic games, encompassing computer, console, and arcade titles, on a game-by-game basis. This extensive catalog includes release details, credits, cover art, player-uploaded screenshots with captions, neutral descriptions, and much more. Impressively, the database boasts a collection of over one million screenshots, meticulously organized by game titles and systems. Additionally, they offer an API that simplifies the process of requesting and retrieving dataset entries and screenshot files. Given these advantages, the Moby Games Database was selected as the primary source for the screenshots used in this research.
@@
To maintain a focused scope for this initial endeavor in video game identification using CNNs, the study exclusively considered home video game consoles. Handheld devices, arcade games, and computer-based titles will be addressed in future research. The selection process involved choosing the top 22 best-selling home video game consoles of all time [23]. These consoles originate from six different manufacturers and exhibit varying quantities of screenshots per game in the database, ranging from none to a few dozen. To ensure that each game has a minimum of five screenshots available for cross-validation purposes, only games meeting this criterion were selected. Table 1 provides an overview of the 22 chosen systems, detailing their total game and screenshot counts, as well as the specific number of games and screenshots selected to satisfy the “at least 5 screenshots” criterion. Figure 1 shows some screenshots from the built dataset.
@@
Xbox Series Microsoft 5 3256 37 44
@@
Table 1: Home Console Video Game Systems, Their Manufacturers, and the Total/Selected Number of Games and Screenshots for the Study. Games with a minimum of five available screenshots in the database were chosen for analysis.
@@
(x) Forza Horizon 5 - Xbox Series
@@ 3 CNN Architectures
In this section, the CNN architectures explored in this study are introduced, along with a description of the additional layers integrated to achieve successful screenshot classification. Table 2 offers an overview of the four architectures under examination, highlighting their input image resolution, the size of their output in the final convolutional layer, the quantity of parameters involved, and citations to their respective references in the literature.
@@
Model Input Image Resolution Output of Last Conv. Layer Parameters Reference
MobileNet $224\times 224$ $7\times 7\times 1024$ $4.3$M Howard et al. [17]
DenseNet169 $224\times 224$ $7\times 7\times 1664$ $14.3$M Huang et al. [18]
EfficientNetB0 $224\times 224$ $7\times 7\times 1280$ $5.3$M Tan and Le [19]
EfficientNetB2 $260\times 260$ $9\times 9\times 1408$ $9.2$M Tan and Le [19]
EfficientNetB3 $300\times 300$ $10\times 10\times 1536$ $12.3$M Tan and Le [19]
Table 2: CNN architectures, some of their characteristics, and their references.
@@
The output from the final convolutional layer of the original CNN is directed into a global average pooling layer. Subsequently, a dropout layer is implemented with a rate of $20\%$ to mitigate overfitting, followed by a large softmax classification layer with up to $1,236$ outputs, which is the amount of game titles in the dataset for the NES console. This proposed architecture is visualized in Figure 2, with $x$ representing the dimensions of the input size (image size), $w$, $y$, and $z$ indicating the dimensions of the CNN output in its ultimate convolutional layer (as detailed in Table 2), and $g$ denoting the output layer’s dimensions, which depend on the number of games from the system being evaluated that are present in the dataset (as indicated in Table 1).
Figure 2: The proposed CNN Transfer Learning architecture.
4 CNN Comparison
@@
This section presents computer simulations that compare CNN models applied to the classification of screenshots from the 22 systems shown in Table 1. All simulations utilized Python and TensorFlow running on three distinct desktop computers equipped with NVIDIA GeForce GPU boards: GTX 970, GTX 1080, and RTX 2060 SUPER, respectively [^1] Access the source code at https://github.com/fbreve/videogame.
For each CNN architecture, image preparation involved resizing to fit the CNN input size and normalizing the range, with no additional preprocessing. Networks commenced with pre-trained weights sourced from the Imagenet dataset [22], which boasts millions of images and hundreds of classes, widely used in transfer learning. These pre-trained weights are available in Tensorflow.
@@
K-Fold Cross Validation, employing $k=5$, was applied universally across all datasets. Training utilized the Adam optimizer [24], initiating with a learning rate of $10^{-3}$ and halving whenever the validation accuracy stagnated for 2 epochs, down to a minimum of $10^{-5}$. Within the training subset, a random $20\%$ of images were allocated to the validation subset, ensuring consistent class proportions through stratification. All models underwent training for up to 50 epochs, with an early stopping criterion to cease training if the validation set loss failed to decrease during the last 10 epochs.
The results, detailed in Table 3, denote averages derived from five different instances of each model, following the Cross Validation approach. They show that DenseNet169 achieved the best accuracy in 14 of the 22 systems. On the other hand, the best average accuracy is attained by EfficientNetB3 ($0.7451$) which is only slightly higher than that achieved by the runner-up DenseNet169 ($0.7446$).
System Mobile- Net Dense- Net169 Efficient- NetB0 Efficient- NetB2 Efficient- NetB3 Average
Atari 2600 0.8925 0.8939 0.8892 0.8920 0.9036 0.8942
NES 0.6550 0.6719 0.6845 0.6876 0.6780 0.6754
Master System 0.7070 0.7452 0.7124 0.7418 0.7429 0.7299
PC Engine 0.7543 0.7989 0.7771 0.7834 0.7890 0.7806
Mega Drive 0.6751 0.7307 0.7326 0.7246 0.7307 0.7188
Super Nintendo 0.6724 0.7247 0.7015 0.7142 0.7081 0.7042
Sega Saturn 0.6973 0.7494 0.6728 0.6841 0.7050 0.7017
PlayStation 0.6908 0.7079 0.6768 0.7227 0.7402 0.7077
Nintendo 64 0.7654 0.8327 0.7383 0.7681 0.7776 0.7764
Dreamcast 0.6521 0.7513 0.6349 0.6735 0.6758 0.6775
PlayStation 2 0.6939 0.7273 0.6836 0.7126 0.7345 0.7104
GameCube 0.7416 0.8233 0.7207 0.7540 0.7498 0.7579
Xbox 0.7804 0.8439 0.7386 0.7929 0.7860 0.7883
Xbox 360 0.7413 0.7859 0.7339 0.7623 0.7768 0.7600
PlayStation 3 0.6218 0.7073 0.6658 0.6890 0.7182 0.6804
Wii 0.7638 0.8337 0.7555 0.7674 0.7840 0.7809
Wii U 0.6594 0.7264 0.6942 0.7006 0.7187 0.6999
PlayStation 4 0.5601 0.6151 0.5975 0.6273 0.6477 0.6096
Xbox One 0.5582 0.6458 0.5791 0.5858 0.6350 0.6008
Nintendo Switch 0.7362 0.8001 0.7511 0.7811 0.7886 0.7714
Xbox Series 0.7643 0.6071 0.9178 0.9714 0.9714 0.8464
PlayStation 5 0.5714 0.6580 0.5863 0.6188 0.6299 0.6129
Average 0.6979 0.7446 0.7111 0.7343 0.7451 0.7266
@@
Table 3: Accuracy achieved by the five different CNN models in each of the $22$ screenshots datasets. Each model is executed five times following the Cross Validation approach. The highest accuracy for each dataset is highlighted in bold.

### v3

@@ Abstract
This paper investigates video game identification through single screenshots, utilizing ten convolutional neural network (CNN) architectures (VGG16, ResNet50, ResNet152, MobileNet, DenseNet169, DenseNet201, EfficientNetB0, EfficientNetB2, EfficientNetB3, and EfficientNetV2S) and three transformers architectures (ViT-B16, ViT-L32, and SwinT) across 22 home console systems, spanning from Atari 2600 to PlayStation 5, totalling 8,796 games and 170,881 screenshots. Except for VGG16, all CNNs outperformed the transformers in this task. Using ImageNet pre-trained weights as initial weights, EfficientNetV2S achieves the highest average accuracy (77.44%) and the highest accuracy in 16 of the 22 systems. DenseNet201 is the best in four systems and EfficientNetB3 is the best in the remaining two systems. Employing alternative initial weights fine-tuned in an arcade screenshots dataset boosts accuracy for EfficientNet architectures, with the EfficientNetV2S reaching a peak accuracy of 77.63% and demonstrating reduced convergence epochs from 26.9 to 24.5 on average. Overall, the combination of optimal architecture and weights attains 78.79% accuracy, primarily led by EfficientNetV2S in 15 systems. These findings underscore the efficacy of CNNs in video game identification through screenshots.
Index Terms:
@@
video game identification, convolutional neural networks, automated game recognition.
I Introduction
@@
Humans possess the remarkable ability to easily recognize their favorite video games or titles they have played frequently from a single screenshot. This proficiency is rooted in the presence of consistent visual elements, including sprites, heads-up displays (HUDs), and distinctive game scenarios. However, extending this capability to the automated identification of random video games from an extensive console library presents a challenge, even for the most dedicated gamers. Therefore, the concept of automatically identifying video games from single screenshots holds significant interest, not only for its technical complexity but also for its vast practical applications.
@@
Automated game identification could offer significant benefits to various sectors within the gaming industry. Video game databases, search engines, and online platforms could gain considerably from this technology. By analyzing user-uploaded screenshots, these platforms can automatically generate metadata, including game titles, release dates, and developer information. Such automation would not only improve the accuracy of their game libraries but also enhance cataloging efficiency. Moreover, online streaming platforms could leverage screenshot recognition to provide real-time information to viewers about the games being played during live streams, enhancing the overall viewer experience. This technology opens doors to further innovation within the gaming ecosystem, potentially influencing game recommendation systems and aiding game-related research.
Most prior research on video game classification has centered around identifying their genre [1, 2, 3, 4]. However, this paper takes a more specific approach, focusing on classifying video game titles based on single screenshots using CNN and transformer models. The hypothesis is that their inherent capacity of automatically extracting relevant features from images is sufficient to identify video game titles from single screenshots in most scenarios, without relying on other features. To begin this research, a dataset encompassing 170,881 screenshots from 8,796 games of 22 popular home console systems was curated. The screenshots were sourced from the Moby Games Database [5]. The proposed dataset spans a wide spectrum of gaming history, ranging from iconic consoles like the ‘Atari 2600’ of the second generation to the cutting-edge ‘PlayStation 5’ and ‘Xbox Series X/S’ of the current generation, carefully selecting the most sold consoles from each generation between them.
To tackle this task, well-established CNN architectures were selected: VGG [6], ResNet [7], MobileNet [8], DenseNet [9], and EfficientNet [10, 11]. These architectures have consistently demonstrated outstanding performance in previous works with different kinds of images [12, 13], making them prime candidates for this game title classification task. Transformers are also being successfully employed in image classification; therefore, two of them were selected for this task: Vision Transformer [14] and Swin Transformer [15]. The initial weights of all those models were first initialized with pre-trained weights from the ImageNet dataset [16], a widely adopted approach in transfer learning [17]. Subsequently, the fine-tuned weights from another dataset of screenshots were employed to enhance both classification accuracy and reduce training times of the best methods for this task. To the best of our knowledge, this marks the first attempt to tackle the challenge of game title classification using CNNs and transformers. It is worth noting that after training a network to identify game titles, incorporating new titles and screenshots into the dataset — whether they are added to the Moby Games Database or provided by the publisher — would necessitate adding additional nodes to the output layer and performing at least a few more training epochs to adjust the new network weights and fine-tune the convolutional layers. By pushing the boundaries of automated video game identification, the aim is to contribute valuable insights to game-related research and practical applications in the ever-evolving gaming industry.
The remainder of this paper is organized as follows. Section II presents related work on video game classification. Section III presents the Moby Games Database and how the dataset was sourced from it. Section IV shows the CNN and transformer architectures employed in this paper. Section V displays the experiments comparing the CNN and transformer architectures in the task of identifying the games from their screenshots, initialized with pre-trained weights from the ImageNet dataset. Section VI demonstrates experiments using weights fine-tuned in another screenshots dataset, comparing the accuracy and training epochs with those obtained with the ImageNet weights and random weights. Section VII presents some limitations of our proposed approach. Finally, the conclusions are drawn in Section VIII.
II Related Work
@@
Most of the video games classification attempts so far aimed at genre classification. Souza et al. [1] pioneering work classified game genre of gameplay videos. Their dataset comprises 700 gameplay videos spanning seven distinct game genres. In their research, they introduced novel descriptors known as Bossa Nova and BinBoost. The experimental outcomes demonstrated the effectiveness of their proposed approach, achieving an accuracy rate of 89.84%.
@@
While the majority of classification endeavors have typically focused on broader categories, there was a unique attempt to classify video games by their titles more than a decade ago. In this pioneering effort, Madani et al. [19] explored several fusion methods using a dataset containing 120,000 gameplay videos, with the objective of identifying game titles. Their approach integrated both audio and visual features to accurately pinpoint these specific game titles, ultimately achieving a F1-score of 0.82. Although their dataset is considerably large, they explored only a small number of games, identifying 30 distinct game titles. To the best of our knowledge, there has not been any other attempt to identify video games by their titles, especially in larger datasets with hundreds or thousands of titles, neither using video nor screenshots.
@@
Recently, Jiang and Zheng [4] devised deep neural networks for the purpose of classifying game genres using either cover images or description text. Their dataset encompassed cover images and description texts sourced from a pool of 50,000 games, which they categorized into 15 distinct genres. In their approach, several pre-trained CNNs were fine-tuned for the cover image classification task. For the classification of description text, they employed Long Short-Term Memory (LSTM) networks and the Universal Sentence Encoder (USE). The image-based model yielded a highest accuracy rate of 31.4% when utilizing ResNet-50. They also achieved significant improvement in accuracy, up to 49.9%, by combining image and text features within a multi-modal model. Both of these previous studies were limited to identifying game genres, and their accuracy was only moderate (below 50%). Therefore, identifying game titles, which are more specific than genres, is still an open challenge.
Video streaming platforms such as YouTube and Twitch have the capability to automatically identify games being played during live streams. However, this functionality appears to be currently limited to a few specific games. In most cases, streamers or uploaders still need to manually label the games they are playing. Unfortunately, these platforms do not openly disclose the methods they use for game identification, whether from existing literature or proprietary development.
III The Dataset
@@
The Moby Games Database [5], as stated on their website, is a project with the primary goal of cataloging comprehensive information about electronic games, encompassing computer, console, and arcade titles, on a game-by-game basis. This extensive catalog includes release details, credits, cover art, player-uploaded screenshots with captions, neutral descriptions, and much more. The database boasts a collection of over one million screenshots, organized by game titles and systems. Additionally, they offer an API that simplifies the process of requesting and retrieving dataset entries and screenshot files. Given these advantages, the Moby Games Database was selected as the source for the screenshots used in this research.
@@
To maintain a focused scope for this initial endeavor in video game identification using CNNs, the study exclusively considered home video game consoles. Handheld devices, arcade games, and computer-based titles will be addressed in future research. The selection process involved choosing the top 22 best-selling home video game consoles of all time [21]. These consoles originate from six different manufacturers and exhibit varying quantities of screenshots per game in the database, ranging from none to a few dozen. In this paper, all experiments employed $k$-fold cross-validation with $k=5$. To ensure that each game has at least one screenshot in each of the five folds, only games with a minimum of five available screenshots were selected. Additionally, during the training phase, it was also assured that each game had screenshots in both the training and validation subsets.
Table I provides an overview of the 22 chosen systems, detailing their total game and screenshot counts, as well as the specific number of games and screenshots selected to satisfy the “at least 5 screenshots” criterion. It is important to note that some of the newer systems, such as the Wii U, Nintendo Switch, Xbox Series X/S, and PlayStation 5, have fewer than a hundred games with available screenshots in Moby Games. Since the last three are current-generation consoles, it is expected that more screenshots will become available over time. However, this current limitation can restrict our analysis. Figure 1 shows some screenshots from the built dataset.
TABLE I: Home Console Video Game Systems, Their Manufacturers, and the Total/Selected Number of Games and Screenshots for the Study. Games with a minimum of five available screenshots in the database were chosen for analysis.
@@
Xbox Series X/S Microsoft 5 3256 37 44
@@
(x) Forza Horizon 5 - Xbox Series X/S
@@ IV CNN and Transformer Architectures
In this section, the CNN and transformer architectures explored in this study are introduced, along with a description of the additional layers integrated to achieve successful screenshot classification. Table II offers an overview of the thirteen architectures under examination, highlighting their input image resolution, the size of their output in the last layer before the classification layer, the quantity of parameters involved, and citations to their respective references in the literature.
TABLE II: CNN and transformer architectures, some of their characteristics, and their references.
@@
Model Input Image Resolution Output of Last Layer Before Classification Parameters Reference
VGG16 $224\times 224$ $7\times 7\times 512$ $138.4$M [6]
ResNet50 $224\times 224$ $7\times 7\times 2048$ $25.6$M [7]
ResNet152 $224\times 224$ $7\times 7\times 2048$ $60.4$M [7]
MobileNet $224\times 224$ $7\times 7\times 1024$ $4.3$M [8]
DenseNet169 $224\times 224$ $7\times 7\times 1664$ $14.3$M [9]
DenseNet201 $224\times 224$ $7\times 7\times 1920$ $20.2$M [9]
EfficientNetB0 $224\times 224$ $7\times 7\times 1280$ $5.3$M [10]
EfficientNetB2 $260\times 260$ $9\times 9\times 1408$ $9.2$M [10]
EfficientNetB3 $300\times 300$ $10\times 10\times 1536$ $12.3$M [10]
EfficientNetV2S $384\times 384$ $12\times 12\times 1280$ $21.6$M [11]
ViT-B16 $224\times 224$ $768$ $8.66$M [14]
ViT-L32 $384\times 384$ $1024$ $306.5$M [14]
SwinT $224\times 224$ $7\times 7\times 768$ $28.3$M [15]
@@
Each of the thirteen network architectures is independently applied to each of the 22 home console datasets, resulting in a total of $286$ variations. The output from the last layer of the original CNN or Swin Transformer (SwinT), before the classification layer, is a 3-D tensor, which is then directed into a global average pooling layer. Subsequently, a dropout layer with a rate of $20\%$ is implemented to mitigate overfitting, followed by a large softmax classification layer with up to 1,236 outputs, corresponding to the number of game titles in the NES console dataset. This proposed architecture is visualized in Figure 2, with $x$ representing the dimensions of the input size (image size), $w$, $y$, and $z$ indicating the dimensions of the CNN or SwinT output in its final layer before classification (as detailed in Table II), and $g$ denoting the output layer’s dimensions, which depend on the number of games from the system being evaluated that are present in the dataset (as indicated in Table I). Note that the Visual Transformer (ViT) models produce 1-D tokens, therefore the pooling layer is not used in their case, and the outputs proceed directly to the dropout layer.
Figure 2: The proposed CNN and SwinT Transfer Learning architecture. The ViT architecture is similar but lacks the pooling layer because its output is one-dimensional.
V Comparison
@@
This section presents experiments that compare the CNN and transformer models applied to the classification of screenshots from each of the 22 systems shown in Table I individually. All experiments utilized Python and TensorFlow running on three distinct desktop computers equipped with four distinct NVIDIA GeForce GPU boards: GTX 970, GTX 1080, RTX 2060 SUPER, and RTX 4060 Ti [^1] Access the source code at https://github.com/fbreve/videogame.
For each CNN and transformer architecture, images were resized to meet the input size requirements of the respective models without maintaining the aspect ratio and were normalized according to each model’s normalization function. No additional preprocessing was performed. The networks were initialized with pre-trained weights from the ImageNet dataset [16], which contains millions of images across hundreds of classes and is commonly used for transfer learning. These pre-trained weights are available with the corresponding CNN or transformer implementations. All experiments utilized Keras implementations of CNNs and transformers [^2] CNNs: https://keras.io/api/applications/, ViT: https://pypi.org/project/vit-keras/, Swin: https://pypi.org/project/tfswin/.
@@
K-Fold Cross Validation, employing $k=5$, was applied universally across all datasets. Training utilized the Adam optimizer [22], initiating with a learning rate of $10^{-3}$ and halving whenever the validation accuracy stagnated for 2 epochs, down to a minimum of $10^{-5}$. In each training stage, from the four folds composing the training subset, a random $20\%$ of images were allocated to the validation subset, ensuring consistent class proportions through stratification. All models underwent training for up to 50 epochs, with an early stopping criterion to cease training if the validation set loss failed to decrease during the last 10 epochs. Other hyperparameters were set to their TensorFlow defaults. These hyperparameters were chosen based on experiments conducted in prior studies using various types of images [12, 13].
The results are detailed in Tables III and IV for CNNs and transformers, respectively. Each result is the average from five different instances of each model, following the Cross Validation approach.
@@
TABLE III: Accuracy achieved by the ten different CNN models in each of the $22$ screenshots datasets. Each model is executed five times following the Cross Validation approach. The highest accuracy for each dataset is highlighted in bold. Standard deviations are shown in parentheses.
System VGG16 ResNet50 ResNet152 MobileNet DenseNet169 DenseNet201 EfficientNetB0 EfficientNetB2 EfficientNetB3 EfficientNetV2S Average
Atari 2600 48,47% (34,87%) 89,25% (0,58%) 87,90% (0,85%) 89,25% (1,16%) 89,39% (1,21%) 89,06% (0,53%) 88,92% (0,98%) 89,20% (1,39%) 90,36% (1,52%) 88,92% (2,74%) 47,26% (44,50%)
NES 50,87% (1,94%) 65,84% (2,30%) 65,65% (0,82%) 65,50% (3,87%) 67,19% (3,65%) 70,04% (1,91%) 68,45% (1,75%) 68,76% (3,27%) 67,80% (4,17%) 72,51% (2,40%) 35,86% (33,11%)
Master System 21,46% (21,46%) 71,85% (0,17%) 66,80% (4,45%) 70,70% (0,77%) 74,52% (0,97%) 73,74% (1,14%) 71,24% (1,00%) 74,18% (0,30%) 74,29% (1,23%) 76,20% (1,19%) 37,00% (35,81%)
PC Engine 30,55% (16,07%) 77,38% (1,26%) 75,36% (1,29%) 75,43% (1,05%) 79,89% (1,39%) 80,19% (0,70%) 77,71% (0,54%) 78,34% (0,83%) 78,90% (0,84%) 80,52% (0,94%) 39,82% (38,07%)
Mega Drive 28,81% (23,31%) 68,43% (2,94%) 68,08% (2,00%) 67,51% (5,76%) 73,07% (2,51%) 73,44% (2,53%) 73,26% (0,43%) 72,46% (3,95%) 73,07% (4,22%) 74,56% (3,22%) 37,74% (36,29%)
Super Nintendo 48,24% (1,70%) 68,23% (2,92%) 68,42% (1,29%) 67,24% (4,92%) 72,47% (3,46%) 74,51% (0,80%) 70,15% (3,00%) 71,42% (3,25%) 70,81% (6,09%) 72,89% (1,72%) 37,23% (34,75%)
Sega Saturn 20,78% (13,98%) 70,04% (1,09%) 67,19% (1,40%) 69,73% (1,42%) 74,94% (2,11%) 75,10% (1,35%) 67,28% (1,53%) 68,41% (1,79%) 70,50% (0,98%) 75,91% (1,11%) 35,80% (34,64%)
PlayStation 27,09% (21,75%) 67,41% (2,33%) 66,10% (2,43%) 69,08% (1,20%) 70,79% (2,78%) 71,19% (2,78%) 67,68% (4,59%) 72,27% (3,02%) 74,02% (1,03%) 74,77% (2,49%) 36,85% (35,32%)
Nintendo 64 23,88% (19,26%) 78,46% (1,56%) 76,10% (1,15%) 76,54% (2,06%) 83,27% (2,07%) 83,67% (1,17%) 73,83% (1,14%) 76,81% (2,65%) 77,76% (0,90%) 82,28% (3,19%) 40,08% (38,73%)
Dreamcast 4,56% (2,25%) 70,47% (1,61%) 67,02% (0,99%) 65,21% (3,81%) 75,13% (1,26%) 75,36% (1,90%) 63,49% (2,62%) 67,35% (3,00%) 67,58% (0,97%) 74,99% (1,83%) 33,74% (33,53%)
PlayStation 2 28,17% (22,71%) 68,87% (0,87%) 66,25% (1,20%) 69,39% (0,83%) 72,73% (1,40%) 74,11% (0,39%) 68,36% (1,13%) 71,26% (2,42%) 73,45% (2,52%) 75,14% (3,29%) 36,68% (35,17%)
GameCube 5,55% (2,92%) 77,49% (1,39%) 75,37% (0,80%) 74,16% (0,69%) 82,33% (0,55%) 82,88% (1,14%) 72,07% (0,86%) 75,40% (3,22%) 74,98% (1,40%) 81,48% (2,30%) 37,16% (36,92%)
Xbox 45,09% (19,18%) 79,80% (1,59%) 78,36% (0,87%) 78,04% (6,20%) 84,39% (1,85%) 84,03% (1,38%) 73,86% (2,62%) 79,29% (1,17%) 78,60% (2,33%) 84,45% (1,32%) 42,14% (39,62%)
Xbox 360 38,72% (17,06%) 74,72% (1,01%) 72,76% (1,15%) 74,13% (0,21%) 78,59% (0,79%) 79,43% (0,72%) 73,39% (0,67%) 76,23% (1,69%) 77,68% (0,86%) 80,25% (1,17%) 39,35% (37,12%)
PlayStation 3 43,80% (1,45%) 63,71% (1,32%) 62,91% (1,60%) 62,18% (5,40%) 70,73% (1,47%) 70,93% (0,90%) 66,58% (1,05%) 68,90% (1,12%) 71,82% (0,92%) 71,98% (3,04%) 35,05% (32,53%)
Wii 20,06% (17,62%) 76,35% (1,54%) 75,82% (1,49%) 76,38% (0,75%) 83,37% (0,90%) 83,05% (0,79%) 75,55% (3,02%) 76,74% (1,90%) 78,40% (1,64%) 83,96% (0,69%) 39,63% (38,55%)
Wii U 14,71% (1,44%) 67,69% (5,09%) 66,06% (3,23%) 65,94% (1,75%) 72,64% (4,73%) 73,03% (2,66%) 69,42% (1,66%) 70,06% (4,44%) 71,87% (2,93%) 74,97% (2,43%) 35,08% (34,39%)
PlayStation 4 36,91% (1,27%) 52,17% (5,04%) 54,38% (2,11%) 56,01% (0,63%) 61,51% (0,41%) 62,19% (0,70%) 59,75% (1,77%) 62,73% (0,98%) 64,77% (1,53%) 63,17% (2,88%) 30,78% (28,70%)
Xbox One 25,99% (4,44%) 61,37% (1,30%) 58,62% (1,03%) 55,82% (1,84%) 64,58% (1,08%) 64,96% (1,33%) 57,91% (1,96%) 58,58% (3,06%) 63,50% (2,92%) 66,25% (2,17%) 31,02% (29,66%)
Nintendo Switch 17,57% (2,95%) 75,63% (2,30%) 70,28% (1,85%) 73,62% (1,09%) 80,01% (1,76%) 79,90% (1,55%) 75,11% (0,60%) 78,11% (1,34%) 78,86% (1,34%) 80,47% (1,70%) 37,80% (36,84%)
Xbox Series X/S 38,21% (6,93%) 43,93% (12,25%) 27,14% (15,75%) 76,43% (16,54%) 60,71% (21,64%) 78,57% (5,98%) 91,78% (6,74%) 97,14% (5,71%) 97,14% (5,71%) 100,00% (0,00%) 41,33% (39,42%)
PlayStation 5 14,95% (3,03%) 59,81% (1,06%) 54,25% (6,38%) 57,14% (1,63%) 65,80% (1,83%) 67,58% (2,22%) 58,63% (3,82%) 61,88% (2,72%) 62,99% (2,57%) 68,10% (2,52%) 30,93% (30,20%)
Average 28,84% (11,71%) 69,50% (2,34%) 66,86% (2,46%) 69,79% (2,89%) 74,46% (2,72%) 75,77% (1,57%) 71,11% (1,98%) 73,43% (2,42%) 74,51% (2,21%) 77,44% (2,02%) 37,20% (35,63%)